# Ring formation from black hole superradiance through repeated particle production on bound orbits

### v1

@@ Abstract
Ultralight bosonic fields around a rotating black hole can extract energy and angular momentum through the superradiant instability and form a dense cloud. We investigate the scenario involving two scalar fields, $\phi$ and $\chi$, with a coupling term $\frac{1}{2}\lambda\phi\chi^{2}$, which is motivated by the multiple-axion framework. The ultralight scalar $\phi$ forms a cloud that efficiently produces $\chi$ particles non-perturbatively via parametric resonance, with a large mass hierarchy, $\mu_{\chi}\gg\mu_{\phi}$. Rather than escaping the system as investigated by previous studies, these $\chi$ particles remain bound, orbiting the black hole. Moreover, the particle production occurs primarily at the peak of the cloud’s profile, allowing $\chi$ particles in quasi-circular orbits to pass repeatedly through resonant regions, leading to sustained amplification. This selective process naturally forms a dense ring of $\chi$ particles, with a mass ratio to the cloud fixed by $(\mu_{\phi}/\mu_{\chi})^{2}$. Our findings reveal a novel mechanism for generating bound-state particles via parametric resonance, which also impacts the evolution of the cloud. This process can be probed through its imprint on binary dynamics and its gravitational-wave signatures.
@@ I Introduction
The interaction between a rotating black hole (BH) and an ultralight bosonic field can trigger a powerful instability known as superradiance Zel’Dovich (1971); Damour et al. (1976); Teukolsky and Press (1974); Press and Teukolsky (1972). (see for a comprehensive review Brito et al. (2015a)). This process extracts rotational energy and angular momentum from the black hole, forming an exponentially growing boson cloud with a hydrogen-like structure, known as a “gravitational atom”. The efficiency of this instability depends on the ratio of the black hole’s gravitational radius to the Compton wavelength of the bosons, which is maximized when the gravitational fine structure constant Baumann et al. (2019a),
@@
This physical insight has transformed astrophysical black holes into laboratories for probing new physics, offering a unique discovery channel for weakly-interacting, ultralight bosons predicted by theories beyond the Standard Model, most notably axion-like particles Arvanitaki et al. (2010); Arvanitaki and Dubovsky (2011). These particles are particularly compelling candidates in this context, as their naturally small mass can be protected by an underlying shift symmetry, broken only by non-perturbative effects Bauer et al. (2021). Furthermore, these particles are well-motivated dark matter candidates Hui et al. (2017). The formation of an axion cloud around a Kerr black hole via superradiance has been a subject of intense research. This includes foundational theoretical investigations Shlapentokh-Rothman (2014); Moschidis (2016); Ficarra et al. (2019); Yang and Huang (2023); Luo and Zhang (2025), numerical studies Dolan (2013); Okawa et al. (2014) and diverse phenomenological studies. The phenomenological explorations cover topics from constraining axion properties with black hole observations Brito et al. (2017); Stott and Marsh (2018); Fernandez et al. (2019); Cheng et al. (2023) to identifying unique gravitational wave signatures Ghosh et al. (2019); Guo et al. (2023); Liu (2024) and assessing the impact of stellar companions on the cloud Baumann et al. (2019b, 2020); Zhang and Yang (2020); Takahashi et al. (2022); Cao and Tang (2023); Tong et al. (2022); Tomaselli et al. (2024). Similar superradiant phenomena for vector and tensor fields have also been extensively explored Pani et al. (2012); Baryakhtar et al. (2017); Cardoso et al. (2018); East and Pretorius (2017); Brito et al. (2020).
@@
The majority of these theoretical studies have focused on the dynamics of a single, non-interacting field. When accounting for self-interactions or couplings to other fields, the superradiant dynamics can be dramatically changed. For instance, self-interactions may lead to scalar emission, level mixing, and the bosenova explosions phenomena Arvanitaki and Dubovsky (2011); Gruzinov (2016); Baryakhtar et al. (2021); Xie and Huang (2025). On the other hand, interactions between the superradiant field and other particle species can be equally consequential, leading to complex dynamics and rich phenomenology Fukuda and Nakayama (2020). Generally, such non-linearity can trigger resonant production of new particles directly from the superradiant condensate. This process drains energy from the cloud by creating a large number of these new particles, thereby suppressing or quenching the superradiant growth Fukuda and Nakayama (2020). Prominent examples of this mechanism include the coupling to photons, which can source powerful, observable bursts of electromagnetic radiation Rosa and Kephart (2018); Ikeda et al. (2019); Sen (2018); Bošković et al. (2019); Spieksma et al. (2023), and the coupling to neutrinos, which can generate significant fluxes of high-energy neutrinos Chen et al. (2025).
@@
In this work, we investigate an interacting two-axion system ($\phi,\chi$) around a rotating black hole, where the lighter superradiant axion-like particle $\phi$ couples to the heavier one $\chi$ via a trilinear coupling, $\frac{1}{2}\lambda\phi\chi^{2}$. Such an interaction drives a powerful, non-perturbative production of $\chi$ particles beyond the kinematic limitations of perturbative decay. We demonstrate a novel ring-formation mechanism that stands in contrast to scenarios where produced particles typically escape the system. Our analysis highlights two key insights. First, particle production commences at the peak of the cloud’s profile, which naturally selects for gravitationally trapped particles on quasi-circular orbits. Second, unlike escaping particles that are amplified only once, these trapped particles repeatedly traverse the resonant regions of the cloud. This leads to sustained amplification, causing their population to grow exponentially and eventually form a stable ring of $\chi$ particles. The existence of multiple, coupled axion fields—an “axiverse”—is well-motivated by fundamental theories like string theory, where a large number of axions with a vast range of masses naturally arise from the compactification of higher-dimensional gauge fields Bauer et al. (2021); Svrcek and Witten (2006); Cicoli et al. (2012); Arvanitaki et al. (2010).
@@
The rest of this paper is organized as follows. In Sec. II, we review the superradiant instability and detail the resonant production of $\chi$ particles from the oscillating $\phi$ cloud. In Sec. III, we analyze the post-production dynamics, employing a WKB approximation to model particle trajectories and demonstrating how sustained amplification leads to the formation of the particle ring. We also discuss the backreaction of the ring on the cloud to determine the final, saturated state of the system. We conclude in Sec. IV. We will use a metric with “mostly plus” signature (,+,+,+)(-,+,+,+)( - , + , + , + ) and work in natural units with $\hbar=c=1$.
@@ II.1 Black hole superradiance
The quasi-bound state solutions have a discrete spectrum of complex eigenfrequencies, $\omega=\omega_{R}+i\omega_{I}$. The superradiance instability arises due to the presence of growing modes for $\phi$, which can be identified as the positive imaginary eigenvalues, i.e., $\omega_{I}>0$. This also implies the superradiance condition
@@
where $\Omega_{\rm H}=\tilde{a}/2r_{+}$ is the angular velocity of the BH’s event horizon, with $\tilde{a}$ the BH dimensionless spin and $r_{+}=GM_{\rm BH}(1+\sqrt{1-\tilde{a}^{2}})$ the horizon radius.
The energy levels represent hydrogen-like features at leading order of $\alpha$ Detweiler (1980); Dolan (2007); Baumann et al. (2019a):
@@
The fastest growing mode $(n\ell m)=(211)$ dominates the scalar field evolution, with the growth rate Detweiler (1980); Dolan (2007); Baumann et al. (2019a):
$\Gamma_{211}=\frac{1}{24}(\tilde{a}-4\alpha)\alpha^{8}\mu\,.$ (5)
The exponential growth of this dominant mode results in the scalar field extracting energy and angular momentum from the black hole, forming a nonspherical, rotating condensate. In the non-relativistic limit, the wavefunction of the $(211)$ state is well-approximated by Chen et al. (2025)
@@
The Newtonian approximation for these wavefunctions is valid in the region where $r\gg r_{g}=GM$. Since the Bohr radius is on the order of $r_{0}=r_{g}/\alpha^{2}$, the approximation is well-justified for small $\alpha$, as the vast majority of the cloud resides far from the black hole. A key feature of this state is that its density profile peaks at a radius $r=2r_{0}$ within the equatorial plane ($\theta=\pi/2$).
When the angular momentum of the BH is extracted to $m\Omega_{\rm H}=\omega_{R}$, i.e., the superradiance condition is no longer met, the scalar cloud saturates. The energy density of the cloud is approximately proportional to the potential energy, $\rho\sim\mu^{2}\phi^{2}$, leading to the total mass of the cloud at saturation Chen et al. (2023),
@@
It has been investigated that the maximal value of $M_{\rm cloud}$ approximates about 10% of the BH mass Brito et al. (2015b); Herdeiro et al. (2022).
@@ II.2 Interacting two-axion superradiant system
We consider a system that contains another scalar $\chi$, which is coupled with $\phi$ through the interaction term $V_{\mathrm{int}}$. The Lagrangian reads
$-\mathcal{L}=\frac{1}{2}(\partial\phi)^{2}+\frac{1}{2}(\partial\chi)^{2}+\frac{1}{2}\mu_{\phi}^{2}\phi^{2}+\frac{1}{2}\mu_{\chi}^{2}\chi^{2}+V_{\mathrm{int}}\,,$ (9)
where $\mu_{\phi}$ and $\mu_{\chi}$ are the masses of the two scalar fields. In this work, we focus on a scenario with a significant mass hierarchy, $\mu_{\phi}\ll\mu_{\chi}$, which can be motivated by the “axiverse” scenarios. [^1] We expect the mass hierarchy would not be extremely large, otherwise we can integrate the heavy field and introduce the effective self-interaction of $\phi$, see Sec.3.4 of Ref. Fukuda and Nakayama (2020) for an example of four-point interaction. The lighter field, $\phi$, satisfies the superradiant condition, meaning its gravitational fine-structure constant $\alpha=GM_{\rm BH}\mu_{\phi}$ is ${\cal O}(0.1)$. The mass hierarchy then implies that the corresponding parameter for the heavier field, $\alpha_{\chi}=GM_{\rm BH}\mu_{\chi}$ is much larger than one, i.e., $\alpha_{\chi}\gg 1$. Consequently, the superradiant instability for the $\chi$ field is strongly suppressed Dolan (2007), and only the $\phi$ field can form a primary cloud via this mechanism.
@@ II.3 Particle production from the trilinear interaction
We now analyze the particle production dynamics induced by the trilinear coupling, $V_{\rm int}=\frac{1}{2}\lambda\phi\chi^{2}$. In this setup, the macroscopic, coherently oscillating $\phi$ cloud acts as a classical background field amplifying fluctuations of the $\chi$ field, leading to exponential particle production. The mechanism of particle production from a coherent scalar field is well-known, most notably from studies of reheating in early-universe cosmology, where a homogeneous inflaton field is responsible for matter production Boyanovsky et al. (1996); Kofman et al. (1997); Bassett et al. (2006). Our scenario introduces crucial distinctions from the typical homogeneous and inflaton background: the superradiant $\phi$ cloud is both spatially inhomogeneous, with a distinct toroidal profile, and its amplitude grows exponentially over time. Initially, when the interaction term is negligible, the amplitude of the $\phi$ field, $\Phi_{0}(t)$, grows exponentially through superradiant instability. Once $\Phi_{0}(t)$ reaches a certain threshold, the oscillation amplitude of $\chi$ field can be enlarged exponentially through parametric resonance.
@@
The $\chi$ field is amplified at the instances when $\omega_{k}^{2}$ is minimal, at which the non-adiabatic condition
|$\absolutevalue{\frac{\dot{\omega}_{k}}{\omega_{k}^{2}}}\gtrsim 1$ (13)
is triggered. In our setup, this occurs on planes where approximately $\varphi=\left(2l+1\right)\pi+\mu_{\phi}t$ ($l\in\mathbb{N}$) and $\phi=-\Phi_{0\star}$. Outside these brief instances of non-adiabaticity, the evolution remains largely adiabatic.
For a fixed amplitude $\Phi_{0\star}$, the range of wavenumbers $k$ that are significantly amplified can be estimated from the stability analysis of the Mathieu equation. In the regime of interest where the pump field is strong ($\lambda\Phi_{0\star}/\mu_{\phi}^{2}\gg 1$), the condition for significant amplification is:
@@
Note that the broad resonance begins with $k^{2}+\mu_{\chi}^{2}\gtrsim\lambda\Phi_{0\star}$, i.e., $\omega_{k}^{2}>0$, due to a minor correction represnent by the second term on the right-hand side, in contrast to tachyonic resonance which requires $\omega_{k}^{2}<0$.
This condition dictates the spatiotemporal onset of particle production. The process becomes efficient only when the cloud’s amplitude $\Phi_{0}$ reaches $\lambda\Phi_{0}\simeq\mu_{\chi}^{2}$. At the threshold, the first modes to be excited are those with the lowest possible $k$ (while remaining consistent with our local approximation, $k\gg\alpha\mu_{\phi}$). Crucially, the profile of the superradiant cloud is not uniform; it has a distinct toroidal profile (as described in Eq. 6) that peaks at a radius of $r=2r_{0}$ in the equatorial plane, $\theta=\pi/2$. Therefore, as the overall cloud grows, this peak toroidal region is the first location in space to reach the resonant threshold. Particle production thus ignites within this specific ring-like zone. As $\Phi_{0}(t)$ continues to grow, the resonance condition is met for progressively larger momentum and across a wider region, opening up the production of more energetic $\chi$ particles.
III Post-production of the $\chi$ particle and formation of the particle ring
In contrast to many scenarios in the literature where particle production is a dissipative process creating escaping relativistic fluxes Rosa and Kephart (2018); Ikeda et al. (2019); Sen (2018); Bošković et al. (2019); Sun and Zhang (2021); Chen et al. (2025), we explore a qualitatively different outcome. As established in the last subsection, the resonance first produces low-momentum particles. We will demonstrate that a select fraction of these low-momentum $\chi$ particles can be gravitationally trapped and confined to stable, quasi-circular orbits within the primary $\phi$ cloud. Unlike all other particles that either escape the system or are quickly accreted, this orbiting population repeatedly traverses the resonant regions. This allows for sustained amplification, a mechanism unique to these bound trajectories, which causes them to become the dominant component of the final state, a critical feature absent in the escaping case. This leads to the formation of a novel toroidal structure of $\chi$ particles in a non-dissipative way for energy redistribution within the superradiant system. To substantiate this picture, we will now derive the classical equations of motion using the WKB analysis within the weak-field limit.

### v2

@@ Abstract
Ultralight bosonic fields around a rotating black hole can extract energy and angular momentum through the superradiant instability and form a dense cloud. We investigate the scenario involving two scalar fields, $\phi$ and $\chi$, with a coupling term $\frac{1}{2}\lambda\phi\chi^{2}$, which is motivated by the multiple-axion framework. The ultralight scalar $\phi$ forms a cloud that efficiently produces $\chi$ particles nonperturbatively via parametric resonance, with a large mass hierarchy, $\mu_{\chi}\gg\mu_{\phi}$. Rather than escaping the system as investigated by previous studies, these $\chi$ particles remain bound, orbiting the black hole. Moreover, the particle production occurs primarily at the peak of the cloud’s profile, allowing $\chi$ particles in quasicircular orbits to pass repeatedly through resonant regions, leading to cumulative amplification. This selective process naturally forms a dense ring of $\chi$ particles, with a mass ratio to the cloud fixed by $(\mu_{\phi}/\mu_{\chi})^{2}$. Our findings reveal a novel mechanism for generating bound-state particles via parametric resonance, which also impacts the evolution of the cloud. This process can be probed through its imprint on binary dynamics and its gravitational-wave signatures.
@@ I Introduction
The interaction between a rotating black hole (BH) and an ultralight bosonic field can trigger a powerful instability known as superradiance zeldo1; Damour:1976kh; Teukolsky:1974yv; Press:1972zz. (see, for a comprehensive review, Brito:2015oca). This process extracts rotational energy and angular momentum from the black hole, forming an exponentially growing boson cloud with a hydrogenlike structure, known as a “gravitational atom.” The efficiency of this instability depends on the ratio of the black hole’s gravitational radius to the Compton wavelength of the bosons, which is maximized when the gravitational fine structure constant Baumann:2019eav,
@@
This physical insight has transformed astrophysical black holes into laboratories for probing new physics, offering a unique discovery channel for weakly interacting, ultralight bosons predicted by theories beyond the Standard Model, most notably axionlike particles Arvanitaki:2009fg; Arvanitaki:2010sy. These particles are particularly compelling candidates in this context, as their naturally small mass can be protected by an underlying shift symmetry, broken only by nonperturbative effects Bauer:2020lowenergy. Furthermore, these particles are well-motivated dark matter candidates Hui:2016ltb. The formation of an axion cloud around a Kerr black hole via superradiance has been a subject of intense research. This includes foundational theoretical investigations Shlapentokh-Rothman:2013ysa; Moschidis:2016wew; Ficarra:2018rfu; Yang:2023vwm; Luo:2024gqo, numerical studies Dolan:2012yt; Okawa:2014nda and diverse phenomenological studies. The phenomenological explorations cover topics from constraining axion properties with black hole observations britoGravitationalWaveSearches2017; Stott:2018opm; Fernandez:2019qbj; chengConstrainsAxionlikeParticle2023; Saha:2022hcd; Bai:2025yxm to identifying unique gravitational wave signatures Ghosh:2018gaw; Guo:2022mpr; liuGravitationalLaserStimulated2024 and assessing the impact of stellar companions on the cloud Baumann:2018vus; baumannGravitationalColliderPhysics2020; Zhang:2019eid; Takahashi:2021yhy; Cao:2023fyv; Tong:2022bbl; Tomaselli:2024dbw. Similar superradiant phenomena for vector and tensor fields have also been extensively explored Pani:2012vp; Baryakhtar:2017ngi; Cardoso:2018tly; East:2017ovw; Brito:2020lup.
@@
The majority of these theoretical studies have focused on the dynamics of a single, noninteracting field. When accounting for self-interactions or couplings to other fields, the superradiant dynamics can be dramatically changed. For instance, self-interactions may lead to scalar emission, level mixing, and the bosenova explosions phenomena Arvanitaki:2010sy; Gruzinov:2016hcq; Baryakhtar:2020gao; Xie:2025npy. On the other hand, interactions between the superradiant field and other particle species can be equally consequential, leading to complex dynamics and rich phenomenology Fukuda:2019ewf. Generally, such nonlinearity can trigger resonant production of new particles directly from the superradiant condensate. This process drains energy from the cloud by creating a large number of these new particles, thereby suppressing or quenching the superradiant growth Fukuda:2019ewf. Prominent examples of this mechanism include the coupling to photons, which can source powerful, observable bursts of electromagnetic radiation rosaStimulatedAxionDecay2018; Ikeda:2018nhb; Sen:2018cjt; boskovicAxionicInstabilitiesNew2019; Spieksma:2023vwl, and the coupling to neutrinos, which can generate significant fluxes of high-energy neutrinos Chen:2023vkq.
@@
In this work, we investigate an interacting two-axion system ($\phi,\chi$) around a rotating black hole, where the lighter superradiant axionlike particle $\phi$ couples to the heavier one $\chi$ via a trilinear coupling, $\frac{1}{2}\lambda\phi\chi^{2}$. Such an interaction drives a powerful, nonperturbative production of $\chi$ particles beyond the kinematic limitations of perturbative decay. We demonstrate a novel ring-formation mechanism that stands in contrast to scenarios where produced particles typically escape the system. Our analysis highlights two key insights. First, particle production commences at the peak of the cloud’s profile, which naturally selects for gravitationally trapped particles on quasicircular orbits. Second, unlike escaping particles that are amplified only once, these trapped particles repeatedly traverse the resonant regions of the cloud. This leads to cumulative amplification, causing their population to grow exponentially and eventually form a stable ring of $\chi$ particles. The existence of multiple, coupled axion fields—an “axiverse”—is well-motivated by fundamental theories like string theory, where a large number of axions with a vast range of masses naturally arise from the compactification of higher-dimensional gauge fields Bauer:2020lowenergy; PeterSvrcek_2006; Cicoli:2012sz; Arvanitaki:2009fg.
@@
The rest of this paper is organized as follows. In Sec. II, we review the superradiant instability and detail the resonant production of $\chi$ particles from the oscillating $\phi$ cloud. In Sec. III, we analyze the postproduction dynamics, employing a WKB approximation to model particle trajectories and demonstrating how cumulative amplification leads to the formation of the particle ring. We also discuss the backreaction of the ring on the cloud to determine the final, saturated state of the system. We conclude in Sec. IV. We will use a metric with “mostly plus” signature (,+,+,+)(-,+,+,+) and work in natural units with $\hbar=c=1$.
@@ II.1 Black hole superradiance
The quasibound state solutions have a discrete spectrum of complex eigenfrequencies, $\omega=\omega_{R}+i\omega_{I}$. The superradiance instability arises due to the presence of growing modes for $\phi$, which can be identified as the positive imaginary eigenvalues, i.e., $\omega_{I}>0$. This also implies the superradiance condition
@@
where $\Omega_{\rm H}=\tilde{a}/2r_{+}$ is the angular velocity of the black hole’s event horizon, with $\tilde{a}$ the black hole’s dimensionless spin and $r_{+}=GM_{\rm BH}(1+\sqrt{1-\tilde{a}^{2}})$ the horizon radius.
The energy levels represent hydrogenlike features at leading order of $\alpha$ detweilerKleinGordonEquationRotating1980; dolanInstabilityMassiveKleinGordon2007a; Baumann:2019eav:
@@
The fastest growing mode $(n\ell m)=(211)$ dominates the scalar field evolution, with the growth rate detweilerKleinGordonEquationRotating1980; dolanInstabilityMassiveKleinGordon2007a; Baumann:2019eav:
$\Gamma_{211}\simeq\frac{1}{48}(\tilde{a}-4\alpha)\alpha^{8}\mu\,.$ (5)
The exponential growth of this dominant mode results in the scalar field extracting energy and angular momentum from the black hole, forming a nonspherical, rotating condensate. In the nonrelativistic limit, the wave function of the $(211)$ state is well approximated by Chen:2023vkq
@@
The Newtonian approximation for these wave functions is valid in the region where $r\gg r_{g}=GM$. Since the Bohr radius is on the order of $r_{0}=r_{g}/\alpha^{2}$, the approximation is well justified for small $\alpha$, as the vast majority of the cloud resides far from the black hole. A key feature of this state is that its density profile peaks at a radius $r=2r_{0}$ within the equatorial plane ($\theta=\pi/2$).
When the angular momentum of the black hole is extracted to $m\Omega_{\rm H}=\omega_{R}$, i.e., the superradiance condition is no longer met, the scalar cloud saturates. The energy density of the cloud is approximately proportional to the potential energy, $\rho\sim\mu^{2}\phi^{2}$, leading to the total mass of the cloud at saturation Chen:2022kzv,
@@
It has been investigated that the maximal value of $M_{\rm cloud}$ approximates about 10% of the black hole mass britoBlackHolesParticle2015; Herdeiro:2021znw.
@@ II.2 Interacting two-axion superradiant system
We consider a system that contains another scalar $\chi$, which is coupled with $\phi$ through the interaction term $V_{\mathrm{int}}$. The Lagrangian reads [^1] In the small-field regime relevant for superradiant growth, expanding the full periodic potential, for instance, $V_{\phi}=\mu_{\phi}^{2}f^{2}[1-\cos\left(\phi/f\right)]$, where $f$ is the axion decay constant, yields the quadratic mass term as well as quartic self-interaction at leading order. Similarly, the effective polynomial interaction term $V_{\mathrm{int}}$ can be derived from a full potential within an effective field theory framework. Although the quartic self-interaction is important, this work focuses on exploring the distinct phenomenology from couplings between different fields. We therefore neglect self-interaction effects to focus on this new dynamic.
$\mathcal{L}=\frac{1}{2}(\partial\phi)^{2}+\frac{1}{2}(\partial\chi)^{2}+\frac{1}{2}\mu_{\phi}^{2}\phi^{2}+\frac{1}{2}\mu_{\chi}^{2}\chi^{2}+V_{\mathrm{int}}\,$, (9)
where $\mu_{\phi}$ and $\mu_{\chi}$ are the masses of the two scalar fields. In this work, we focus on a scenario with a significant mass hierarchy, $\mu_{\phi}\ll\mu_{\chi}$, which can be motivated by the “axiverse” scenarios. [^2] We expect the mass hierarchy would not be extremely large, otherwise we can integrate the heavy field and introduce the effective self-interaction of $\phi$, see Sec.3.4 of Ref. Fukuda:2019ewf for an example of four-point interaction. The lighter field, $\phi$, satisfies the superradiant condition, meaning its gravitational fine-structure constant $\alpha=GM_{\rm BH}\mu_{\phi}$ is ${\cal O}(0.1)$. The mass hierarchy then implies that the corresponding parameter for the heavier field, $\alpha_{\chi}=GM_{\rm BH}\mu_{\chi}$ is much larger than one, i.e., $\alpha_{\chi}\gg 1$. Consequently, the superradiant instability for the $\chi$ field is strongly suppressed dolanInstabilityMassiveKleinGordon2007a, and only the $\phi$ field can form a primary cloud via this mechanism.
@@ II.3 Particle production from the trilinear interaction
We now analyze the particle production dynamics induced by the trilinear coupling, $V_{\rm int}=\frac{1}{2}\lambda\phi\chi^{2}$. In this setup, the macroscopic, coherently oscillating $\phi$ cloud acts as a classical background field amplifying fluctuations of the $\chi$ field, leading to exponential particle production. The mechanism of particle production from a coherent scalar field is well known, most notably from studies of reheating in early-universe cosmology, where a homogeneous inflaton field is responsible for matter production Boyanovsky:1996sq; Kofman:1997yn; Bassett:2005xm. Our scenario introduces crucial distinctions from the typical homogeneous and inflaton background: the superradiant $\phi$ cloud is both spatially inhomogeneous, with a distinct toroidal profile, and its amplitude grows exponentially over time. Initially, when the interaction term is negligible, the amplitude of the $\phi$ field, $\Phi_{0}(t)$, grows exponentially through superradiant instability. Once $\Phi_{0}(t)$ reaches a certain threshold, the oscillation amplitude of $\chi$ field can be enlarged exponentially through parametric resonance.
@@
The $\chi$ field is amplified at the instances when $\omega_{k}^{2}$ is minimal, at which the nonadiabatic condition
[…]
is triggered. In our setup, this occurs on planes where approximately $\varphi=\left(2l+1\right)\pi+\mu_{\phi}t$ ($l\in\mathbb{N}$) and $\phi=-\Phi_{0\star}$. Outside these brief instances of nonadiabaticity, the evolution remains largely adiabatic.
For a fixed amplitude $\Phi_{0\star}$, the range of wave numbers $k$ that are significantly amplified can be estimated from the stability analysis of the Mathieu equation. In the regime of interest where the pump field is strong ($\lambda\Phi_{0\star}/\mu_{\phi}^{2}\gg 1$), the condition for significant amplification is
@@
Note that the broad resonance begins with $k^{2}+\mu_{\chi}^{2}\gtrsim\lambda\Phi_{0\star}$, i.e., $\omega_{k}^{2}>0$, due to a minor correction represnented by the second term on the right-hand side, in contrast to tachyonic resonance which requires $\omega_{k}^{2}<0$.
This condition dictates the spatiotemporal onset of particle production. The process becomes efficient only when the cloud’s amplitude $\Phi_{0}$ reaches $\lambda\Phi_{0}\simeq\mu_{\chi}^{2}$. At the threshold, the first modes to be excited are those with the lowest possible $k$ (while remaining consistent with our local approximation, $k\gg\alpha\mu_{\phi}$). Crucially, the profile of the superradiant cloud is not uniform; it has a distinct toroidal profile (as described in Eq. (6)) that peaks at a radius of $r=2r_{0}$ in the equatorial plane, $\theta=\pi/2$. Therefore, as the overall cloud grows, this peak toroidal region is the first location in space to reach the resonant threshold. Particle production thus ignites within this specific ringlike zone. As $\Phi_{0}(t)$ continues to grow, the resonance condition is met for progressively larger momentum and across a wider region, opening up the production of more energetic $\chi$ particles.
III Postproduction of the $\chi$ particle and formation of the particle ring
In contrast to many scenarios in the literature where particle production is a dissipative process creating escaping relativistic fluxes rosaStimulatedAxionDecay2018; Ikeda:2018nhb; Sen:2018cjt; boskovicAxionicInstabilitiesNew2019; sunFastGravitationalWave2021; Chen:2023vkq, we explore a qualitatively different outcome. As established in the last subsection, the resonance first produces low-momentum particles. We will demonstrate that a select fraction of these low-momentum $\chi$ particles can be gravitationally trapped and confined to stable, quasicircular orbits within the primary $\phi$ cloud. Unlike all other particles that either escape the system or are quickly accreted, this orbiting population repeatedly traverses the resonant regions. This allows for cumulative amplification, a mechanism unique to these bound trajectories, which causes them to become the dominant component of the final state, a critical feature absent in the escaping case. This leads to the formation of a novel toroidal structure of $\chi$ particles in a nondissipative way for energy redistribution within the superradiant system. To substantiate this picture, we will now derive the classical equations of motion using the WKB analysis within the weak-field limit.